OpenStudy (anonymous):

i will fan and medal Kenny solved the equation t plus two-thirds equals one-sixth as follows: Line one shows: t plus two-thirds minus two-thirds equals one-sixth minus two-thirds. Line two shows: t equals one-sixth minus two-thirds. Line three shows: t equals one-sixth minus four-sixths. Line four shows: t equals one-half. Which statement best describes his error? A. He subtracted the wrong fraction on both sides of the equal sign. B. He subtracted when he should have added two-thirds to both sides of the equal sign. C. He did not rename the fractions correctly when

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

give me a sec im figuring it out

OpenStudy (anonymous):

what does C say it looks chopped off

OpenStudy (mathstudent55):

@12345! This is a math problem. Does your keyboard have number keys? Instead of spelling out the numbers, it's much easier to understand a math problem if you see numbers instead of the numbers spelled out. It would also take you much less time to write out the problem.

OpenStudy (mathstudent55):

Original equation: \(t + \dfrac{2}{3} = \dfrac{1}{6}\) \(1. ~~~t + \dfrac{2}{3} - \dfrac{2}{3}= \dfrac{1}{6} - \dfrac{2}{3}\) \(2. ~~~t = \dfrac{1}{6} - \dfrac{2}{3}\) \(3. ~~~t = \dfrac{1}{6} - \dfrac{4}{6}\) \(4. ~~~t = \dfrac{1}{2}\)

OpenStudy (mathstudent55):

Let's go over lines 1 through 4. To solve this equation, you should first subtract 2/3 from each side. That is what line 1 shows, so line 1 is correct.

OpenStudy (mathstudent55):

On the left side, 2/3 - 2/3 = 0, leaving only t, so line 2 is correct.

OpenStudy (mathstudent55):

In line 3, you see that the fraction 2/3 was changed to 4/6 to have the common denominator, 6, needed for the subtraction. Since 2/3 = 4/6, line 3 is also correct.

OpenStudy (mathstudent55):

From line 3 to line 4, you need to subtract fractions on the right side. 1/6 - 4/6 = -3/6 3/6 reduces to 1/2, but the fraction is not 3/6, it's -3/6, so the answer is t = -1/2, not t = 1/2. The error was in step 4.